OpenStudy (babynini):

Simplifying: tan(theta) + cos (-theta) + tan(-theta) how do I get rid of negative thetas?

OpenStudy (babynini):

tan(Θ)+cos(−Θ)+tan(−Θ) (sinΘ/cosΘ)+cos(−Θ)+(sin/cos)(−Θ)

zepdrix (zepdrix):

Hey Nini :) Cosine is an even function meaning, \(\Large\rm \cos(-\theta)=\cos(\theta)\) Sine and Tangent are odd functions therefore, \(\Large\rm \tan(-\theta)=-\tan(\theta)\) and \(\Large\rm \sin(-\theta)=-\sin(\theta)\)

zepdrix (zepdrix):

Do you understand how to apply these properties to your problem? No need to convert to sines and cosines ^^

OpenStudy (babynini):

Hello:) haha you made my night by saying hi, thank you! Um, yes I think so. The original equation would be: \[\tan(\Theta) +\cos(\Theta) - \tan(\Theta)\] which would come out to: cos(Theta) ?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (babynini):

yayay!